OpenStudy (anonymous):

Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No Solution

jimthompson5910 (jim_thompson5910):

|x - 2| + 10 = 12 |x - 2| = 12-10 |x - 2| = 2 x - 2 = 2 or x - 2 = -2 x = 2+2 or x = -2+2 x = 4 or x = 0

OpenStudy (anonymous):

@jim Solve for x: |x + 2| + 16 = 14 x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution

OpenStudy (anonymous):

x=4

jimthompson5910 (jim_thompson5910):

|x + 2| + 16 = 14 |x + 2| = 14-16 |x + 2| = -2 Since the absolute value of any number is nonnegative, this means that the last equation is not possible So there are no solutions.

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

just two more @jim Solve for x: −5|x + 1| = 10 x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution

OpenStudy (anonymous):

one more actually

jimthompson5910 (jim_thompson5910):

−5|x + 1| = 10 |x + 1| = 10/(-5) |x + 1| = -2 So there are no solutions (same reasoning as given above)

OpenStudy (anonymous):

LAST ONE @jim Solve for x: |3x + 2| = 14 x = 4 and x = −4 x = 4 and x = -16/3 x = and x = -16/3 No Solutions

jimthompson5910 (jim_thompson5910):

|3x + 2| = 14 3x + 2 = 14 or 3x + 2 = -14 3x = 14-2 or 3x = -14-2 3x = 12 or 3x = -16 x = 12/3 or x = -16/3 x = 4 or x = -16/3

OpenStudy (anonymous):

@jim im so sorry but those were the wrong question could you please solve these problems with just the answer IM SO SORRY Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No Solution Question 2 (Multiple Choice Worth 2 points) Solve for x: |x| = 4 x = 0 and x = 4 x = 4 and x = −4 x = 4 x = −4 Question 3 (Multiple Choice Worth 2 points) Solve for x: −5|x + 1| = 10 x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution Question 4 (Multiple Choice Worth 2 points) Solve for x: |3x + 2| = 2 x = 0 and x = -4/3 x = and x = -4/3 x = 0 No Solutions Question 5 (Multiple Choice Worth 2 points) Solve for x: |x + 2| + 16 = 14 x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution

jimthompson5910 (jim_thompson5910):

# 1 x = 9 or x = -15 ------------------------------------------------------- # 2 x = 4 or x = −4 ------------------------------------------------------- # 3 No Solution ------------------------------------------------------- # 4 x = 0 or x = -4/3 ------------------------------------------------------- # 5 No Solution

OpenStudy (anonymous):

so the answer is no solution